OpenStudy (anonymous):

Solve the separable differential equation dy/dx=(x-10)e^(-y), given the initial conditions y(2)=ln(2), to find y(0)

OpenStudy (anonymous):

wouldn't you jut use implicit differentiation and then substitute?

OpenStudy (anonymous):

ok first you find a diff of y=ln(x), and replace in the equation, is most easy, sorry i dont have a good english

OpenStudy (anonymous):

\[dy = (x-10)dx*e^-(x)\] \[e^y*dy = (x-10)dx\]

OpenStudy (anonymous):

then integrate both sides

OpenStudy (anonymous):

I'm not sure I understand. In my answer I used partial derivatives, but I'm not sure how I got the answer when I did it previously...

OpenStudy (anonymous):

fxx(x,y)=6x and fyy(x,y)=2 and fxy(x,y)=-6, and I got D=6x(2)-(-6)^2=12x-36 which is the answer... but I don't know where I got 6x(2)-(-6)^2 from previously.

OpenStudy (anonymous):

^Sorry, the above post was meant for a different question.

OpenStudy (anonymous):

Frayar22, do you mean differentiate y=ln(x) ... y=1/x? I don't get it :(

OpenStudy (anonymous):

sorry , i have a mistake sorry forgive all i say (sorry for my bad englsih)